A deep sea diver is at a depth of 125m below sea level. If the density of sea water is 1025kgm-³ where g=10Nkg-¹ determine the pressure on the diver due to the sea water. If atmospheric pressure is 101000Pa, determine the total pressure on the diver
1
Expert's answer
2019-02-20T13:53:19-0500

We can determine the pressure on the diver due to the sea water from the formula:


"P = \rho_{sea \thinspace water} gh,"


here, "P" is the pressure on the diver due to the sea water, "\rho_{sea \thinspace water}" is the density of the sea water, "g" is the acceleration due to gravity and "h" is the depth.

Then, we get:


"P_{sea \thinspace water} = 1025 \dfrac{kg}{m^3} \cdot 10 \dfrac{N}{kg} \cdot 125 m = 1.28 \cdot 10^6 Pa."


We can find the total pressure on the diver from the formula:


"P_{total} = P_{atm} + P_{sea \thinspace water} = 1.01 \cdot 10^5 Pa + 1.28 \cdot 10^6 Pa = 1.38 \cdot 10^6 Pa."

Answer:

"P_{sea \thinspace water} = 1.28 \cdot 10^6 Pa."

"P_{total} = 1.38 \cdot 10^6 Pa."